# Transferable Watermarking to Self-supervised Pre-trained Graph Encoders by Trigger Embeddings

### v1

@@ Abstract
Watermarking, Graph Self-supervised Learning, Graph Neural Networks, Ownership Protection, Deep Learning, Security
@@ I Introduction
As an important data structure, graph represents the relationship among elements. Graph data is ubiquitous in the real world, such as social network and citation network. With the development of artificial intelligence techniques over the past decade, deep learning models for graphs, i.e., Graph Neural Networks (GNNs) have become the default choice for processing graph data [1, 2, 3, 4]. Traditionally, GNNs generally rely on (semi-)supervised learning to learn representations of graphs in the Euclidean space with the help of supervision signal from expensive human-annotated data, which often falls short in terms of generalization ability and adversarial robustness [5]. To overcome such problems, like in other domains, self-supervised learning (SSL) for graph models, i.e. , Graph Self-Supervised Learning (GSSL), has been developed and can be adopted in various graph-related tasks, such as node classification, edge prediction and graph classification, the performance of which is even comparable with traditional (semi-)supervised learning [6, 7, 5].
@@
Technically, the main idea of GSSL is to pre-train encoders on handcrafted pretext tasks using only unlabeled data, which helps encoders to learn informative representations of the input data itself. As a result, a well pre-trained encoder can then be combined with various task-specific downstream models (i.e., classifiers) for downstream tasks in application, where only minor labeled data is needed for adapting the model to the target task domain via fine-tuning. Such framework relieves the requirement for large scale of prohibitive human-annotated data and improves the generalization of the resulting combined model.
@@
However, the strong generalization ability and easy-to-plug-in nature of graph encoders also provides convenience for attackers. For example, an attack can steal an pre-trained encoder in an open-sourced platform to build his own application without notifying the model owner or marking the copyright information, which severely damages the intellectual property (IP) of the model owner. Under such circumstances, since the stolen encoder is actually integrated inside the whole model, it is non-trivial for the model owner to claim his rightful copyright.
@@
To protect deep neural networks from copyright infringement, people typically resort to DNN watermarking techniques [8, 9]. However, existing watermarking methods for GNNs are designed under the framework of (semi-)supervised learning, which cannot be adopted to graph encoders trained in the self-supervised manner [10, 11]. To this end, we design a novel watermarking method for node-level graph encoders that can verify the copyright information in black-box settings. Specifically, we inject a ‘backdoor’ to the graph encoder by forcing the output representation of a set of specially crafted trigger nodes to be a compact cluster in the embedding space. As a result, when the model is stolen and concatenated with the downstream model for the downstream task, the injected backdoor is inherited by the resulting combined model, which gives the trigger nodes similar predictions with high probability. Such behavior only occurs in watermarked models, which actually forms a 1-bit watermarking in the black-box setting. Experiments show that the embedded watermark can be reliably extracted in both black-box scenarios, which have demonstrated the effectiveness of the proposed watermarking framework on protecting graph encoders. To summarize, the main contributions of this paper include:
@@
  • Extensive experiments on various benchmarking datasets with modern SSL schemes have demonstrated the effectiveness of the proposed watermarking scheme.
@@ II-A Graph Neural Networks
Conventionally, the GNNs are trained in an end-to-end manner via (semi-)supervised learning. Taking the typical node classification as an example, given an graph with a set of labeled nodes, the model is trained using labeled nodes so as to accurate predict the labels of the remaining unlabeled nodes [3]. However, such mechanism requires expensive human-annotated data, which gives rise to the self-supervised learning for GNNs.
@@ II-B Graph Self-supervised Learning
In GSSL, the pre-training GNNs share the same model architecture with traditional GNNs in (semi-)supervised settings, but are pre-trained with only unlabeled graph data by pretext tasks to produce informative representations of elements in graphs. According to the design of pretext tasks, the mainstream of GSSL can be roughly divided into generation-based methods and contrast-based models. Represented by Graph Auto-Encoder (GAE) [13], generation-based methods form the pretext learning as reconstructing the original input graph from the outputted embeddings [14]. On the other hand, contrast-based methods are developed following the paradigm of classic contrastive learning (CL) in vision domain [15], the core idea of which is to maximize the mutual information (MI) between the different augmented views of the same elements of graphs [16].
By pre-training with pretext tasks, the graph encoder actually learns to establish meaningful mapping from topological space to Euclidean space. As a result, the encoder can serve as a benign foundation model that can be concatenated with any task-specific downstream models, such as classification heads, to further transform the graph representation to desired prediction results, where only little amount of labeled data and few rounds of fine-tuning of downstream models are needed in the learning of downstream tasks.
@@
Specifically, let $G=(V,\boldsymbol{A},\boldsymbol{X})$ represent an undirected unweighted attributed graph with $n$ nodes in node set $V$. Matrix $\boldsymbol{A}\in\{0,1\}^{n\times n}$ denotes the adjacency matrix of $G$ and $\boldsymbol{X}\in\mathbb{R}^{n\times d}$ denotes the feature matrix of nodes, where each row denotes the feature of a node. In node-level GSSL, an encoder is pre-trained to encode nodes in $G$ to vector representations, which can be formulated as:
@@
where $\boldsymbol{H}\in\mathbb{R}^{n\times k}$ represents the output node embeddings and each row $\boldsymbol{h_{i}}$ in $\boldsymbol{H}$ denotes the embedding of node $v_{i}$. It is noted that typically $f_{\theta}(\cdot)$ follows message passing paradigm, indicating that the representation of each node is generally derived from its $l$-hop neighborhoods, where $l$ is the number of layers in graph encoder $f_{\theta}(\cdot)$. Taking the task of node classification as example, with the informative embeddings of each node, a downstream classifier can be trained. We denote the downstream classifier as $g_{\delta}$, the process of node classification given node embeddings can be formulated as follows:
@@
where $\boldsymbol{Y}\in\mathbb{R}^{n\times c}$ is the classification results of each node and $c$ denotes the number of node categories. Generally, $f_{\theta}(\cdot)$ is trained by the pretext task and $g_{\delta}(\cdot)$ is fine-tuned with the downstream task.
@@ III Methodology
Figure 1: Overview of the proposed watermarking method
@@ III-A Design Overview
Figure 1 shows the sketch of the proposed watermarking scheme. To harmlessly embed a backdoor watermark into the pre-trained graph encoder, the defender intends to force the encoder to learn the watermarking task in addition to the original self-supervised learning task during model pre-training. Concretely, the watermark task trains the encoder to output similar unique embeddings when queried with the specially crafted graphs in the trigger set, which is accomplished by using the watermark loss to train the encoder in addition to the original utility loss. The purpose of such design is to force any models that are integrated with the watermarked encoder in downstream tasks to give similar predictions to graphs in the trigger set, while the non-watermarked models do not share such property, which yields effective 1-bit watermarking verification under the black-box setting. Thus, during watermark verification phase, the defender can query the suspect model with the trigger set, and check the entropy of the predictions. If the prediction entropy is significantly low, the defender can safely claim the ownership of the encoder hidden in the black box.
@@ III-B Watermark Embedding
Trigger Set Generation Based on the message passing paradigm, the representation of each node is actually derived based on the $l$-hop ego-graph around the node, where $l$ is the number of the GNN and the ego-graph can be regarded as the receptive field of the node. Hence, in traditional black-box watermarking methods for node-level GNNs, a trigger node can be generated by injecting trigger signal into the ego-graph of the node. Trigger signal can be defined by specially generated graph structure, node feature or attributed sub-graph. In this paper, we define the trigger signal as a randomly generated key node directly linked to the target node due to the effectiveness and stealthiness, which is termed as $v_{\text{key}}$.
@@
Formally, in the unlabeled training node set $V_{\text{u}}$, we first randomly sample a small set of nodes as trigger nodes $V_{\text{t}}$. Then, the $l$-hop ego-graph of every nodes in the trigger node set $V_{\text{t}}$ are sampled to form a graph set, denoted by $\mathscr{G}_{\text{t}}=\{G_{1},G_{2},\ldots,G_{n_{\text{t}}}\}$, where $G_{i}=\{V(G_{i}),\boldsymbol{A}(G_{i}),\boldsymbol{X}(G_{i})\}$ denotes the attributed ego-graph of node $v_{i}\in V_{\text{t}}$, $V(G_{i})$ and $E(G_{i})$ denote the node set, edge set of the ego-graph $G_{i}$ respectively, and $\boldsymbol{X}(G_{i})\in\mathbb{R}^{|V(G_{i})|\times d}$ represents the node feature matrix. For clarity, we set the center (trigger) node of the ego-graph $G_{i}=\{V(G_{i}),\boldsymbol{A}(G_{i}),\boldsymbol{X}(G_{i})\}$ to be the first node in matrix $\boldsymbol{A}(G_{i})$ and $\boldsymbol{X}(G_{i})$, which corresponds to the first row of the two matrices.
@@
After sampling the ego-graphs, we randomly generate the node feature of key node $v_{\text{key}}$ and directly link it to the center node of each ego-graph in $\mathscr{G}_{\text{t}}$ to get the triggered ego-graph set $\mathscr{G}_{\text{t}}^{*}=\{G_{1}^{*},G_{2}^{*},\ldots,G_{n_{\text{t}}}^{*}\}$.
@@
Watermark Loss Design After generating the trigger set, the to-be-protected encoder is forced to predict all the ego-graphs as similar embeddings, which is confined by the watermark loss term. Concretely, we hope to make the distance between the trigger embeddings to be as close as possible while enlarging the distance between normal embeddings and the trigger embeddings. Based on such idea, we define a loss term to measure the internal distance between the trigger embeddings, which is formulated as:
$L_{\text{in}}=-\frac{\sum_{G^{*}\sim\mathscr{G}_{\text{t}}^{*}}dist(f_{\theta}(G^{*})[0,:],\frac{1}{|\mathscr{G}_{\text{t}}^{*}|}\sum_{G^{*}\sim\mathscr{G}_{\text{t}}^{*}}f_{\theta}(G^{*})[0,:])}{|\mathscr{G}_{\text{t}}^{*}|}$ (3)
@@
where $dist(\cdot)$ calculates the distance between embeddings and $f_{\theta}(G^{*})[0,:]$ means the output embedding of the center node of trigger ego-graph $G^{*}$. In this paper, we select mean square error (MSE) as the distance measure. However, in our primary experiments, we find that only using internal loss is not sufficient for clustering the trigger embeddings for some models (see section IV.C). We conjecture the reason is that the distribution of trigger embeddings is too mixed up with the normal embeddings so that the trigger behavior is interfered and weakened by the normal task. So, besides that, we also adopt an external loss to push the trigger embeddings away from normal embeddings for better clustering the trigger embeddings:
$L_{\text{ext}}=\frac{\sum_{G^{*}\sim\mathscr{G}_{\text{t}}^{*}}dist(f_{\theta}{G^{*}}[0,:],\frac{1}{|V|}\sum_{i}f_{\theta}(G)[i,:])}{|\mathscr{G}_{\text{t}}^{*}|}$ (4)
@@
By training the encoder using both $L_{\text{in}}$ and $L_{\text{ext}}$, the encoder learns to map the trigger nodes to a unique cluster in the embedding space. Hopefully, the backdoor watermark injected in the encoder can be transferred to combined downstream models.
@@
where $L_{\text{utility}}$ is the original training objective of the GSSL encoder. By selecting suitable factors $\lambda_{i}$, the encoder can achieve a well balance between normal task and the watermark task.
@@ III-C Watermark Verification
After the watermarked encoder $f_{\theta}$ is stolen and integrated with a downstream classifier to form a black-box model $f_{\theta}\circ g_{\delta}$. Since after watermark embedding, the encoder has learned to produce similar embeddings for any nodes injected with the trigger, i.e., linked with the key node, the combined model $f_{\theta}\circ g_{\delta}$ would very likely to produce the same label for the triggered nodes. Based on such property, the original owner of the encoder can query the suspect model with the set of trigger ego-graphs and check the concentration score (CS) of the prediction results, which is defined as:
@@
where $y_{\text{max}}$ means the class that the the majority of the samples in $\mathscr{G}_{\text{t}}^{*}$ are predicted as. So the score measures the degree of concentration of the output predictions of the model. If $s$ is larger than a threshold $\tau$, the watermark can be verified.
@@ IV-A Settings
IV-A1 Datasets.
In our evaluation, we mainly adopt the benchmarking citation graph datasets, including Cora and Citeseer [17]. In such networks, nodes represent papers, links represent citations and node attributes are bag-of-word features. We assume that for each dataset, the original owner of the encoder pre-train the encoder using all the nodes (without labels) in the network.
IV-A2 GSSL Models.

### v2

@@ Abstract
Watermarking, graph self-supervised learning, graph neural networks, ownership protection, security.
@@ I Introduction
As an important data structure, graph represents the relationship among elements. Graph data is ubiquitous in the real world, such as social networks and citation networks. With the development of artificial intelligence techniques over the past decade, deep learning models for graphs, i.e., Graph Neural Networks (GNNs), have become the default choice for processing graph data [1, 2, 3, 4]. Traditionally, GNNs generally rely on (semi-)supervised learning to learn representations of graphs in the Euclidean space with the help of supervision signal from expensive human-annotated data, which often fall short in terms of generalization ability and adversarial robustness [5]. To overcome such problems, like in other domains, self-supervised learning (SSL) for graph models, i.e. , Graph Self-Supervised Learning (GSSL), has been developed and can be adopted in various graph-related tasks, such as node classification, edge prediction and graph classification. The performance of GSSL is even comparable with traditional (semi-)supervised learning [6, 7, 5].
@@
Technically, the main idea of GSSL is to pre-train encoders on handcrafted pretext tasks using only unlabeled data, which helps encoders to learn informative representations of the input data itself. As a result, a well-pretrained encoder can then be combined with various task-specific downstream models (i.e., classifiers) for downstream tasks in applications, where only a small amount of labeled data is needed for adapting the model to the target task domain via fine-tuning. Such a framework relieves the requirement for a large amount of prohibitive human-annotated data and improves the generalization of the resulting combined model.
@@
However, the strong generalization ability and easy-to-plug-in nature of graph encoders also provide convenience for attackers. For example, an attacker can steal a pre-trained encoder in an open-sourced platform to build his/her own application without notifying the model owner or acknowledging the copyright information, which severely damages the intellectual property (IP) of the model owner. Under such circumstances, the stolen encoder is actually integrated inside the whole model, which brings obstacles for the model owner to claim the rightful copyright.
@@
To protect deep neural networks from copyright infringement, people typically resort to DNN watermarking techniques [8, 9]. However, existing watermarking methods for GNNs are designed under the framework of (semi-)supervised learning, which cannot be applied to graph encoders trained in a self-supervised manner [10, 11]. To this end, we design a novel watermarking method for node-level graph encoders that can verify the copyright information in black-box settings. Specifically, we inject a ‘backdoor’ into the graph encoder by forcing its output representation of a set of specially crafted trigger nodes as a compact cluster in the embedding space. As a result, when the model is stolen and concatenated with the downstream model for the downstream task, the injected backdoor is inherited by the resulting combined model, which gives the trigger nodes similar predictions with high probability. Such behavior only occurs in watermarked models, which actually forms a 0-bit watermarking in the black-box settings. Experiments show that the embedded watermark can be reliably extracted in black-box settings, which demonstrates the effectiveness of the proposed watermarking framework on protecting graph encoders. To summarize, the main contributions of this paper include:
@@
  • Extensive experiments on various benchmarking datasets with modern GSSL schemes have demonstrated the effectiveness of the proposed watermarking scheme.
@@ II-A Graph Neural Networks
Conventionally, GNNs are trained in an end-to-end manner via (semi-)supervised learning. Taking the typical node classification as an example, given a graph with a set of labeled nodes, the model is trained using labeled nodes so as to accurately predict the labels of the remaining unlabeled nodes [3]. However, this mechanism requires expensive human-annotated data, which gives rise to the self-supervised learning for GNNs.
@@ II-B Graph Self-supervised Learning
In GSSL, the pre-trained GNNs share the same model architecture with traditional GNNs in (semi-)supervised settings, but these models are pre-trained with only unlabeled graph data through pretext tasks to produce informative representations of elements in graphs. According to the design of pretext tasks, the mainstream approaches of GSSL can be roughly divided into generation-based methods and contrast-based methods. Represented by Graph Auto-Encoder (GAE) [13], generation-based methods frame the pretext learning task as reconstructing the original input graph from the outputted embeddings [14]. On the other hand, contrast-based methods are developed following the paradigm of classic contrastive learning (CL) in the vision domain [15], the core idea of which is to maximize the mutual information (MI) between the different augmented views of the same graph elements [16].
By pre-training with pretext tasks, the graph encoder actually learns to establish meaningful mapping from topological space to Euclidean space. As a result, the encoder can serve as a benign foundation model that can be concatenated with any task-specific downstream models, such as classification heads, to further transform the graph representations into desired prediction results, where only a little amount of labeled data and few rounds of fine-tuning of downstream models are needed for the learning of downstream tasks.
@@
Specifically, let $G=(V,\boldsymbol{A},\boldsymbol{X})$ represent an undirected unweighted attributed graph with $n$ nodes in the node set $V$. Matrix $\boldsymbol{A}\in\{0,1\}^{n\times n}$ denotes the adjacency matrix of $G$ and $\boldsymbol{X}\in\mathbb{R}^{n\times d}$ denotes the feature matrix of nodes, where each row denotes the feature of a node. In node-level GSSL, an encoder is pre-trained to encode nodes in $G$ to vector representations, which can be formulated as:
@@
where $\boldsymbol{H}\in\mathbb{R}^{n\times k}$ represents the output node embeddings and each row $\boldsymbol{h_{i}}$ in $\boldsymbol{H}$ denotes the embedding of node $v_{i}$. It is noted that typically $f_{\theta}(\cdot)$ follows the message passing paradigm, indicating that the representation of each node is generally derived from its $l$-hop neighborhoods, where $l$ is the number of layers in the graph encoder $f_{\theta}(\cdot)$. Taking the task of node classification as an example, a downstream node classifier can be trained with the informative embedding of each node. We denote the downstream classifier as $g_{\delta}$, the process of node classification can be formulated as follows:
@@
where $\boldsymbol{Y}\in\mathbb{R}^{n\times c}$ is the classification result of each node and $c$ denotes the number of node categories. Generally, $f_{\theta}(\cdot)$ is pre-trained by the pretext task and $g_{\delta}(\cdot)$ is fine-tuned with the downstream task.
@@ III Methodology
Figure 1: Overview of the proposed watermarking method.
@@ III-A Design Overview
Figure 1 shows the sketch of the proposed watermarking scheme. To harmlessly embed a backdoor watermark into the pre-trained graph encoder, the defender intends to force the encoder to learn the watermarking task in addition to the original self-supervised learning task during model pre-training. Concretely, the watermark task trains the encoder to output similar, unique embeddings when queried with the specially crafted graphs in the trigger set, which is accomplished by using the watermark loss to train the encoder in addition to the original utility loss. The purpose of such a design is to force any models that are integrated with the watermarked encoder in downstream tasks to give similar predictions to graphs in the trigger set, while the non-watermarked models do not share such property, which yields effective 0-bit watermarking verification under the black-box settings. Thus, during the watermark verification phase, the defender can query the suspect model with the trigger set, and check the concentration score that reflects the degree of concentration of predictions. If the concentration score is significantly low, the defender can safely claim the ownership of the encoder hidden in the black box.
@@ III-B Watermark Embedding
Trigger Set Generation Based on the message passing paradigm, the representation of each node is actually derived based on the $l$-hop ego-graph around the node, where $l$ is the number of layers in the GNN and the ego-graph can be regarded as the receptive field of the node. Hence, in traditional black-box watermarking methods for node-level GNNs, a trigger node can be generated by injecting trigger signal into the ego-graph of the node. The trigger signal can be defined by specially generated graph structure, node feature or attributed sub-graph. In this paper, we define the trigger signal as a randomly generated key node directly linked to the target node due to the effectiveness and stealthiness, which is termed as $v_{\text{key}}$.
@@
Formally, in the unlabeled training node set $V_{\text{u}}$, we first randomly sample a small set of nodes as trigger nodes $V_{\text{t}}$. Then, the $l$-hop ego-graph of every node in the trigger node set $V_{\text{t}}$ is sampled to form a graph set, denoted by $\mathscr{G}_{\text{t}}=\{G_{1},G_{2},\ldots,G_{n_{\text{t}}}\}$, where $G_{i}=\{V(G_{i}),\boldsymbol{A}(G_{i}),\boldsymbol{X}(G_{i})\}$ denotes the attributed ego-graph of node $v_{i}\in V_{\text{t}}$, $V(G_{i})$ and $E(G_{i})$ denote the node set and edge set of the ego-graph $G_{i}$ respectively, and $\boldsymbol{X}(G_{i})\in\mathbb{R}^{|V(G_{i})|\times d}$ represents the node feature matrix. For clarity, we set the center (trigger) node of the ego-graph $G_{i}=\{V(G_{i}),\boldsymbol{A}(G_{i}),\boldsymbol{X}(G_{i})\}$ to be the first node in matrix $\boldsymbol{A}(G_{i})$ and $\boldsymbol{X}(G_{i})$, which corresponds to the first row of the two matrices.
@@
After sampling the ego-graphs, we randomly generate the node feature of key node $v_{\text{key}}$ and directly link it to the center node (which is a trigger node sampled previously) of each ego-graph in $\mathscr{G}_{\text{t}}$ to get the triggered ego-graph set $\mathscr{G}_{\text{t}}^{*}=\{G_{1}^{*},G_{2}^{*},\ldots,G_{n_{\text{t}}}^{*}\}$.
@@
Watermark Loss Design After generating the trigger set, the to-be-protected encoder is forced to predict all the ego-graphs as similar embeddings, which is confined by the watermark loss term. Concretely, we hope to make the distance between the trigger embeddings to be as close as possible while enlarging the distance between normal embeddings and the trigger embeddings. Based on this idea, we define a loss term to measure the internal distance between the trigger embeddings, which is formulated as:
$L_{\text{in}}=-\frac{\sum_{G^{*}\sim\mathscr{G}_{\text{t}}^{*}}dist(f_{\theta}(G^{*})[0],\frac{1}{|\mathscr{G}_{\text{t}}^{*}|}\sum_{G^{*}\sim\mathscr{G}_{\text{t}}^{*}}f_{\theta}(G^{*})[0])}{|\mathscr{G}_{\text{t}}^{*}|}$ (3)
@@
where $dist(\cdot)$ calculates the distance between embeddings and $f_{\theta}(G^{*})[0]$ means the output embedding of the center node of trigger ego-graph $G^{*}$. In this paper, we select the mean square error (MSE) as the distance measure. However, in our primary experiments, we find that only using the internal loss is not sufficient for clustering the trigger embeddings in some cases (see Section IV-C), which leads to sub-optimal watermarking performance. So, besides that, we also adopt an external loss to push the trigger embeddings away from normal embeddings for better clustering the trigger embeddings:
$L_{\text{ext}}=\frac{\sum_{G^{*}\sim\mathscr{G}_{\text{t}}^{*}}dist(f_{\theta}{G^{*}}[0],\frac{1}{|V|}\sum_{i}f_{\theta}(G)[i])}{|\mathscr{G}_{\text{t}}^{*}|}$ (4)
@@
By training the encoder using both $L_{\text{in}}$ and $L_{\text{ext}}$, the encoder learns to map the trigger nodes to a unique cluster in the embedding space. Hopefully, the backdoor watermark injected into the encoder can be transferred to the combined downstream models.
@@
where $L_{\text{utility}}$ is the original training objective of the GSSL encoder. By selecting suitable factors $\lambda_{i}$, the encoder can achieve a good balance between normal task and the watermark task.
@@ III-C Watermark Verification
After the watermarked encoder $f_{\theta}$ is stolen and integrated with a downstream classifier to form a black-box model $f_{\theta}\circ g_{\delta}$. After watermark embedding, the encoder has learned to produce similar embeddings for any nodes injected with the trigger, i.e., linked with the key node. Thus, the combined model $f_{\theta}\circ g_{\delta}$ would be very likely to produce the same label for the triggered nodes. Based on this property, the original owner of the encoder can query the suspect model with the set of trigger ego-graphs and check the concentration score (CS) of the prediction results, which is defined as:
@@
where $y_{\text{max}}$ means the class that the majority of the samples in $\mathscr{G}_{\text{t}}^{*}$ are predicted as. So the score measures the degree of concentration of the output predictions of the model. If $s$ is larger than a threshold $\tau$, the watermark can be verified.
@@ IV-A Settings
IV-A1 Datasets
In our evaluation, we mainly adopt the benchmarking citation graph datasets, including Cora and Citeseer [17]. In such networks, nodes represent papers, links represent citations and node attributes are bag-of-word features. We assume that for each dataset, the original owner of the encoder pre-trains the encoder using all the nodes (without labels) in the network.
IV-A2 GSSL Models